# Text-only adaptation in LLM-based ASR through text denoising

### v2

@@ Abstract
Adapting automatic speech recognition (ASR) systems based on large language models (LLMs) to new domains using text-only data is a significant yet underexplored challenge. Standard fine-tuning of the LLM on target-domain text often disrupts the critical alignment between speech and text modalities learned by the projector, degrading performance. We introduce a novel text-only adaptation method that emulates the audio projection task by treating it as a text denoising task. Our approach thus trains the LLM to recover clean transcripts from noisy inputs. This process effectively adapts the model to a target domain while preserving cross-modal alignment. Our solution is lightweight, requiring no architectural changes or additional parameters. Extensive evaluation on two datasets demonstrates up to 22.1% relative improvement, outperforming recent state-of-the-art text-only adaptation methods.
@@ 1 Introduction
footnotetext: Corresponding authors: {sergio.burdisso, esau.villatoro}@idiap.ch
@@
In this context, LLM-based ASR systems have emerged as a practical and computationally efficient alternative, focusing on transcription by leveraging a fixed, manually defined prompt during both training and inference [16, 22, 26, 12, 25, 7, 20, 24, 1]. This fixed-prompt setup ensures consistency between the training objective and inference behavior, making it well-suited for applications where high-accuracy transcription is the primary goal. A key advantage of LLM-based ASR is the ease of combining strong pretrained speech encoders with powerful LLMs through a learnable projection layer, thereby leveraging advances from pretraining in both speech and text modalities. This modular design enables scalable, high-performance transcription without the need for costly instruction tuning. Intuitively, the projection layer learns to map speech representations into the text embedding space of the LLM (i.e., learns a speech-to-text alignment). Once projected, the resulting representation can be interpreted by the LLM as a noisy text, which the model reconstructs into a clean transcription through its inherent denoising capability.
@@
To address these challenges, we propose a novel text-only adaptation strategy that fine-tunes the LLM within an LLM-based ASR architecture by means of formulating the problem as a denoising task. Our contributions are as follows: (i) We reformulate text-only adaptation as a denoising task, training the LLM to reconstruct inputs that mimic the outputs of a speech projector in LLM-based ASR architectures. (ii) We propose a lightweight training approach that simply consists of a multi-view noise-driven batching strategy, not requiring additional learnable parameters. (iii) We present a thorough evaluation across two datasets, achieving up to 22.1% relative improvement, surpassing the state-of-the-art. (iv) We release our source code for reproducibility. [^1] https://github.com/idiap/llm-asr-text-adaptation
Fig. 1: Batch composition used for fine-tuning the LLM during text-only adaptation to a target domain. Here, $\sigma$ and $\boldsymbol{\tau}$ denote the proportions of the batch that are drawn from the source domain ($\mathcal{D}_{src}$) and the target domain ($\mathcal{D}_{tgt}$), respectively.
@@ 2 Method
LLM-based ASR systems consist of three main components: (i) a pretrained speech encoder (e.g., Whisper [17], Hubert [10], or WavLM [2], etc.) that extracts frame-level acoustic representations, (ii) a learnable projector $sp(\cdot)$ that maps these representations into the LLM input embedding space (e.g., pooling-based [4],CNN-based [5, 18], or linear-based [16, 8, 12]), [^2] Also known as “speech projector” in the literature, with some referring to it as a “speech adapter” or “connector”. and (iii) a pre-trained LLM (e.g., Llama [6], Vicuna [3] ) that acts as a decoder, and generates final transcripts.
Prior work has shown that even a simple projector $sp(\cdot)$ (two linear layers with a nonlinearity) combined with a frozen LLM is sufficient to obtain strong transcription performance [16, 12, 25]. This suggests that the projector learns to convert speech into a sequence of soft tokens that approximate entries in the LLM vocabulary. For example, the projector may map the audio utterance “yes that would be” into embeddings resembling “mmy Z YesssS S SGS that Will B be S S”, as reported in prior work [12, 25]. This illustrates how the projector outputs resemble a noisy transcript rather than raw speech features. As a result, the LLM is forced to interpret these inputs as a corrupted or noisy version of the transcript, which it then recovers.
@@ 2.1 Task Formulation
Let $\mathcal{D}_{src}=\{(a_{i},t_{i})\}$ denote the source-domain dataset with paired audio $a_{i}$ and transcript $t_{i}$, and let $\mathcal{D}_{tgt}=\{t_{j}\}$ denote the target-domain dataset with transcripts only. [^3] It is worth mentioning that in all our experiments, the text-only data consists of (ground truth) transcripts from conversational speech, rather than arbitrary text such as that found on web pages. While standard LLM-based ASR training requires large numbers of $(a,t)$ pairs to achieve high recognition performance, we enable text-only adaptation by introducing a noise function $noise(\cdot)$. This function takes transcripts as input and emulates the audio projection task by adding synthetic noise that approximates the outputs of a trained projector in an LLM-based architecture.
@@ 2.2 Batch Construction for Text Denoising Adaptation
Directly fine-tuning the LLM with $(noise(t),t)$ pairs from $\mathcal{D}_{tgt}$ leads to catastrophic forgetting, i.e., the alignment between the speech encoder and the LLM degrades, and the projector’s mapping becomes ineffective, negatively impacting the model’s performance. A similar observation has also been recently reported in  [7]. To mitigate this, we propose a more effective batch composition strategy for fine-tuning the LLM. Specifically, each batch is constructed as a mixture of examples from both source and target domains (see Figure 1). Let $\sigma_{a}$, $\sigma_{ta}$, $\sigma_{t}$, and $\tau_{t}$ denote the proportions (relative shares) of the following components in each batch:
@@
  • $\sigma_{ta}$: $(noise_{a}(t),t)$ pairs where $(a,t)\in\mathcal{D}_{src}$ and $noise_{a}(t)$ is obtained by projecting $a$ through the model’s projector (i.e. $sp(a)$) and mapping the resulting embeddings to their nearest tokens in the LLM vocabulary. This approximates an optimal noise function, as it corresponds to projector-induced text noise.
@@ 4 Experimental Setup
Prompt(i, t) = <|start_header_id|>user
<|end_header_id|>Transcribe speech to text. Speech:{i}<|eot_id|><|start_header_id|>
assistant<|end_header_id|>{t}
where {t} is the ground truth transcript and {i} is the input that will change with the type of (input, output) pair in the batch composition, in particular:
  • Prompt($sp(a)$, $t$) for $(a,t)$
@@
  • Prompt($noise(t)$, $t$) for $(noise(t),t)$,
@@
$\bullet$ Adapted model (audio): This experiment reflects the best-case scenario, i.e., when audio-text pairs are available for fine-tuning the LLM for the target domain. Specifically, for adapting the LLM, we fine-tuned on the target partition, using audio-text pairs, for four epochs, using LoRA applied to the self-attention query and value projection layers with rank 8, alpha 32, learning rate $1\!\times\!10^{-4}$, and $1000$ warm-up steps.
$\bullet$ Text-only adaptation (ours): For these experiments, we implement $noise(t)$ as a two-step process: (1) random character substitution followed by (2) random character duplication. In step (1), we use nlpaug [^7] https://nlpaug.readthedocs.io/en/latest/augmenter/char/random.html to select $15\%$ of the words and replace $30\%$ of the characters within those words with random symbols, with a minimum of 1 and a maximum of 10 character edits per utterance. These values match the nlpaug defaults, except that we reduce the fraction of edited words from $30\%$ to $15\%$, which yielded better performance in preliminary experiments. In step (2), each character has a $10\%$ chance of being repeated; when selected, it is duplicated 1 to 3 times with equal probability. This last step is designed to emulate the duplication patterns observed in the optimal noise (i.e. $noise_{a}(t)$).
@@
$\bullet$ Ma et al. [15]: the text-only adaptation is performed in two stages: (1) first, freeze the model and learn $k$ trainable embeddings $e_{i}$ inserted where the audio would go in the prompt —i.e. the input of the LLM is Prompt($e_{1}\dots e_{k}$, $t$); (2) fine-tune the LLM using those $k$ embeddings instead of audio, i.e., batches are composed only of $(e_{1}\dots e_{k},t)$ pairs. In our experiments, we use $k=30$ as in the original work.
@@
Finally, all experiments are evaluated in terms of word error rate (WER), and we report the relative improvement ($\Delta$) over the corresponding base model.
@@ 5 Experimental Results
We evaluate our proposed text-only adaptation method in three scenarios of increasing difficulty: (1) in-domain adaptation, (2) out-of-domain adaptation, and (3) cross-domain adaptation. The main results for each of these experiments are reported in Table 2, Table 3 and Table 4, respectively.
@@ 5.1 Ablation experiments
Table 6: Ablation of different batch item type for $\sigma_{t}$ and $\tau$. The original noise is replaced by different strategies. $\Delta$ is reported over the base model of Table 2. Values in %.
Strategy Batch Item Type Banking Insurance
@@
Noise Prompt($noise(t)$, $t$) $10.11$ 22.1 $8.71$ 17.9
Echoing Prompt($t$, $t$) $10.31$ 20.6 $9.09$ 14.3
Empty Prompt($\emptyset$, $t$) $10.56$ 18.6 $9.01$ 15.1
@@ 6 Conclusions
We have presented a novel text-only adaptation approach for LLM-based ASR architectures. Our method alternates source audio, projector-induced noise, and synthetic noisy text during fine-tuning, enabling the model to retain audio understanding, learn text denoising, and acquire target-domain knowledge. Experiments show consistent gains across domains, outperforming prior text-only adaptation methods. Future work will explore improved noise functions that better approximate projection outputs and evaluate optimal $\tau$ values under text-rich, real-world conditions.
Acknowledgments This work was supported by an Idiap Research Institute and Uniphore collaboration project. Part of this work was also supported by EU Horizon 2020 project ELOQUENCE.

### v3

@@ Abstract
Adapting large language model (LLM)-based automatic speech recognition (ASR) systems to new domains using text-only data is a significant yet underexplored challenge. Standard fine-tuning of the LLM on the target domain text often disrupts the critical alignment between the speech and text modality learned by the projector, degrading performance. We introduce a novel text-only adaptation method that frames this process as a text denoising task. Our approach trains the LLM to recover clean transcripts from noisy inputs. This process effectively adapts the model to a target domain while preserving cross-modal alignment. Our solution is lightweight, requiring no architectural changes or additional parameters. Extensive evaluation on two datasets demonstrates up to 22.1% relative improvement, outperforming recent state-of-the-art text-only adaptation methods.
@@ 1 Introduction
footnotetext: Performed experimentation and data preparation. Conceived the original idea and helped plan the experiments. Involved in planning and supervised the work, writing the manuscript with support from and . Provided critical feedback and helped shape the research.
@@
In this context, LLM-based ASR systems have emerged as a practical and computationally efficient alternative, focusing on transcription by leveraging a fixed, manually defined prompt during both training and inference [3, 4, 12, 13, 14, 15, 16, 17]. This fixed-prompt setup ensures consistency between the training objective and inference behavior, making it well-suited for applications where high-accuracy transcription is the primary goal. A key advantage of LLM-based ASR is the ease of combining strong pretrained speech encoders with powerful LLMs through a learnable projection layer, thereby leveraging advances from pretraining in both speech and text modalities. This modular design enables scalable, high-performance transcription without the need for costly instruction tuning. Intuitively, the projection layer learns to map speech representations into the text embedding space of the LLM (i.e., learns a speech-to-text alignment). Once projected, the resulting representation can be interpreted by the LLM as a noisy text, which the model reconstructs into a clean transcription through its inherent denoising capability.
@@
To address these challenges, we propose a novel text-only adaptation strategy that fine-tunes the LLM within an LLM-based ASR architecture by means of formulating the problem as a denoising task. Our contributions are as follows: (i) We reformulate text-only adaptation as a denoising task, training the LLM to reconstruct inputs that mimic the outputs of a speech projector in LLM-based ASR architectures. (ii) We propose a lightweight training approach that simply consists of a multi-view noise-driven batching strategy, not requiring additional learnable parameters. (iii) We present a thorough evaluation across two datasets, achieving up to 22.1% relative improvement, surpassing the state-of-the-art.
Fig. 1: Batch composition used for fine-tuning the LLM during text-only adaptation to a target domain. Here, $\sigma$ and $\boldsymbol{\tau}$ denote the proportions of the batch that are drawn from the source domain ($\mathcal{D}{src}$) and the target domain ($\mathcal{D}{tgt}$), respectively.
@@ 2 Method
LLM-based ASR systems consist of three main components: (i) a pretrained speech encoder (e.g., Whisper [19], Hubert [20], or WavLM [21], etc.) that extracts frame-level acoustic representations, (ii) a learnable projector (e.g., pooling-based [2],CNN-based [11, 22], or linear-based [3, 23, 13]) that maps these representations into the LLM input embedding space, [^1] Also known as “speech projector” in the literature, with some referring to it as a “speech adapter” or “connector”. and (iii) a pre-trained LLM (e.g., Llama [24], Vicuna [25] ) that acts as a decoder, and generates final transcripts.
Prior work has shown that even a simple projector (two linear layers with a nonlinearity) combined with a frozen LLM is sufficient to obtain strong transcription performance [3, 13, 14]. This suggests that the projector learns to convert speech into a sequence of soft tokens that approximate entries in the LLM vocabulary. For example, the projector may map the audio utterance “yes that would be” into embeddings resembling “mmy Z YesssS S SGS that Will B be S S”, as reported in prior work [13, 14]. This illustrates how the projector outputs resemble a noisy transcript rather than raw speech features. As a result, the LLM is forced to interpret these inputs as a corrupted or noisy version of the transcript, which it then recovers.
@@ 2.1 Task Formulation
Let $\mathcal{D}_{src}=\{(a_{i},t_{i})\}$ denote the source-domain dataset with paired audio $a_{i}$ and transcript $t_{i}$, and let $\mathcal{D}_{tgt}=\{t_{j}\}$ denote the target-domain dataset with transcripts only. [^2] It is worth mentioning that in all our experiments, the text-only data consists of (ground truth) transcripts from conversational speech, rather than arbitrary text such as that found on web pages. While standard LLM-based ASR training requires large numbers of $(a,t)$ pairs to achieve high recognition performance, we enable text-only adaptation by introducing a noise function $noise(\cdot)$. This function takes as input transcripts and generates perturbed variants that approximate the outputs of a trained projector in the LLM-based architecture.
@@ 2.2 Batch Construction for Text Denoising Adaptation
Directly fine-tuning the LLM with $(noise(t),t)$ pairs from $\mathcal{D}_{tgt}$ leads to catastrophic forgetting, i.e., the alignment between the speech encoder and the LLM degrades, and the projector’s mapping becomes ineffective, negatively impacting the model’s performance. A similar observation has also been recently reported in  [15]. To mitigate this, we propose a more effective training batch composition for fine-tuning the LLM. Specifically, each batch is constructed as a mixture of examples from both source and target domains (see Figure 1). Let $\sigma_{a}$, $\sigma_{ta}$, $\sigma_{t}$, and $\tau_{t}$ denote the proportions (relative shares) of the following components in each batch:
@@
  • $\sigma_{ta}$: $(noise_{a}(t),t)$ pairs where $(a,t)\in\mathcal{D}_{src}$ and $noise_{a}(t)$ is obtained by projecting $a$ through the model’s projector and mapping the resulting embeddings to their nearest tokens in the LLM vocabulary. This approximates an optimal noise function, as it corresponds to projector-induced text noise.
@@ 4 Experimental Setup
Prompt(input) = <|start_header_id|>user
<|end_header_id|>Transcribe speech to text. Speech:{input}<|eot_id|><|start_header_id|>
assistant<|end_header_id|>{transcript}
Notice that the value of {input} will change with the type of (input, output) pair in the batch composition, in particular:
  • Prompt($a$) for $(a,t)$
@@
  • Prompt($noise(t)$) for $(noise(t),t)$,
where {transcript} is always the ground truth transcript $t$.
@@
$\bullet$ Adapted model (audio): This experiment reflects the best-case scenario, i.e., when audio-text pairs are available for fine-tuning the LLM for the target domain. Specifically, for adapting the LLM, we fine-tuned on the target partition, using audio-text pairs, for four epochs, using LoRA applied to the self-attention query and value projection layers with rank 8, learning rate $1\!\times\!10^{-4}$, and $1000$ warm-up steps.
$\bullet$ Text-only adaptation (ours): For these experiments, we implement noise(text) as a two-step process: (1) random character substitution followed by (2) random character duplication. In step (1), we use nlpaug [^6] https://nlpaug.readthedocs.io/en/latest/augmenter/char/random.html to select $15\%$ of the words and replace $30\%$ of the characters within those words with random symbols, with a minimum of 1 and a maximum of 10 character edits per utterance. These values match the nlpaug defaults, except that we reduce the fraction of edited words from $30\%$ to $15\%$, which yielded better performance in preliminary experiments. In step (2), each character has a $10\%$ chance of being repeated; when selected, it is duplicated 1 to 3 times with equal probability. This step is designed to emulate the duplication patterns observed in noise(audio).
@@
$\bullet$ Ma et al. [18]: the text-only adaptation is performed in two stages: (1) first, freeze the model and learn $k$ trainable embeddings $e_{i}$ inserted where the audio would go in the prompt (i.e. the input of the LLM is Prompt($e_{1}\dots e_{k}$)); (2) fine-tune the LLM using those $k$ embeddings instead of audio, i.e., batches are composed only of $(e_{1}\dots e_{k},t)$ pairs. In our experiments, we use $k=30$ as in the original work.
@@
Finally, all experiments are evaluated in terms of word error rate (WER), and report the relative improvement ($\Delta$) over the corresponding base model.
@@ 5 Experimental Results
We evaluate our proposed text-only adaptation method in three scenarios of increasing difficulty: (1) in-domain adaptation, (2) out-of-domain adaptation, (3) cross-domain adaptation. The main results for each of these experiments are reported in Table 2, Table 3 and Table 4, respectively.
@@ 5.1 Ablation experiments
Table 6: Batch composition with respect to the LLM input format for $\sigma_{t}$ and $\tau$. The original noise is replaced by different strategies. $\Delta$ is reported over the base model of Table 2. Values in %.
Strategy Input Banking Insurance
@@
Noise Prompt($noise(t)$) $10.11$ 22.1 $8.71$ 17.9
Echo Prompt($t$) $10.31$ 20.6 $9.09$ 14.3
Empty prompt Prompt() $10.56$ 18.6 $9.01$ 15.1
@@ 6 Conclusions
This paper presented a novel text-only adaptation approach for LLM-based ASR architectures that mitigates catastrophic forgetting when fine-tuning with textual data alone. Our method integrates multiple components during fine-tuning—alternating between source audio, projector-induced text noise, and synthetic noisy text—allowing the model to simultaneously (i) preserve its ability to interpret audio-conditioned inputs and (ii) learn to denoise text, while also acquiring target-domain knowledge. Experiments demonstrate consistent improvements across multiple domains and datasets, surpassing prior state-of-the-art text-only adaptation strategies. As future work, we plan to explore more sophisticated noise functions that better approximate the projection layer outputs and to conduct a thorough evaluation of the $\tau$ value to assess performance under text-rich, real-world conditions.